# Opportunities and Challenges of Large Language Models for Low-Resource Languages in Humanities Research

### v2

@@ 1.1.2 Demand for Low-Resource Languages in Humanities Research and the Insufficiencies in Existing Studies
The critical significance of low-resource languages notwithstanding, prevailing research methodologies encounter appreciable challenges when engaging with these languages. A principal limitation is the paucity of textual data. Numerous low-resource languages lack extensive written corpora, thereby complicating the execution of empirical studies and the development of computational models. Furthermore, the absence of specialized computational tools specifically adapted to these languages aggravates the issue [57]. Existing tools and resources are frequently designed for high-resource languages, thereby proving ineffective or inefficient for application to low-resource languages. This technological disparity is further heightened by the inadequate availability of essential linguistic resources such as dictionaries, grammars, and annotated corpora, which are imperative for both descriptive and computational linguistic research.
@@ 1.3 Research Objectives and Contributions
2 Foundational Framework for LLMs in Low-Resource Language Research
2.1 Definitions and Types of Low-Resource Languages
2.1.1 Dialects
The challenges in processing dialects include:
2.1.2 Ancient Languages
2.1.3 Endangered Languages
Endangered languages are at risk of extinction, often spoken by small, isolated communities. Many indigenous languages in regions such as Australia, Africa, and the Americas face dwindling numbers of speakers, limited geographical reach, and minimal documentation.
The challenges in endangered language processing include:
To address these challenges, large language models can leverage transfer learning, zero-shot learning, and unsupervised methods to process limited data effectively. Furthermore, community-driven initiatives can play a pivotal role in collecting and annotating linguistic data.
2.2 Scarcity and Complexity of Corpora in Low-Resource Languages
The available datasets for low-resource languages often suffer from poor quality and limited scope, making them suboptimal for robust model training [80]. Inconsistencies within the datasets and the absence of comprehensive linguistic coverage further restrict their utility. The problem is exacerbated by the underrepresentation of these languages in AI research, as commercial interests have historically focused on high-resource languages. This imbalance has resulted in technological advancements that predominantly benefit languages with larger speaker bases and more substantial economic influence.
2.3 Role of Large Language Models in Addressing Challenges
Large language models offer a transformative approach to addressing the challenges posed by low-resource languages. By leveraging advanced techniques such as few-shot learning, meta-learning, and unsupervised pre-training, these models can mitigate the effects of data scarcity and complexity. Some promising strategies include:
  • Transfer learning: Pre-training on high-resource languages and fine-tuning on related low-resource languages.
  • Synthetic data generation: Using generative models to create artificial corpora that replicate linguistic features of low-resource languages.
  • Collaborative annotation: Engaging native speakers and local communities in the creation of datasets to ensure linguistic authenticity.
2.4 Techniques Supporting Low-Resource Languages
2.4.1 Transfer Learning
Key benefits of transfer learning include:
2.4.2 Cross-Language Pretraining
Cross-language pretraining is another powerful method, exemplified by multilingual models like mBERT, XLM, and XLM-R [27]. These models are trained on multilingual corpora, enabling them to learn shared representations across languages.
Advantages of cross-language pretraining include:
  • Shared representations: Multilingual models encode universal linguistic features, facilitating cross-linguistic knowledge transfer.
  • Scalability: These models can support a wide range of languages without the need for extensive language-specific fine-tuning.
  • Support for zero-shot and few-shot learning: Cross-language pretraining enables models to perform well on unseen low-resource languages by leveraging knowledge from related languages.
For instance, mBERT has demonstrated effectiveness in tasks like cross-language question answering and translation, where shared multilingual embeddings provide a robust foundation for low-resource language processing.
2.4.3 Multi-Task Learning
Applications of multi-task learning for low-resource languages include:
2.4.4 Data Augmentation
Data augmentation addresses the problem of insufficient training data by generating additional examples [79]. This technique enriches the dataset, providing the model with more diverse and representative samples for training.
Common data augmentation methods include:
  • Synthetic data generation: Machine translation can create parallel corpora by translating high-resource language datasets into low-resource languages.
  • Paraphrasing: Generating alternative expressions for existing sentences to increase linguistic diversity.
  • Noise injection: Adding controlled noise to data to simulate real-world variability and improve robustness.
For example, back-translation has proven effective in tasks like machine translation and text classification, where augmenting datasets with machine-generated examples significantly boosts model performance.
2.4.5 Multi-Modal Integration
Advantages of multi-modal integration include:
2.5 Adaptation and Gaps Between LLMs and Low-Resource Languages
In recent years, LLMs have made significant progress and demonstrated immense potential in the field of humanities research. These models can process vast amounts of text, uncover hidden information, and provide new perspectives and methods for tasks such as ancient text interpretation and cultural analysis. This capability is particularly beneficial for humanities research, especially for low-resource languages, where LLMs can serve as a critical tool to break research bottlenecks, enhancing the understanding, analysis, and preservation of these languages. However, there are still limitations in the adaptation of current LLMs to low-resource languages that need to be thoroughly examined.
Availability of Corpora The available corpora for low-resource languages are diverse but fraught with issues. On one hand, the sources include digitization projects of historical documents, language archives from specific regions, and limited academic research collections. However, these sources are scattered and significantly smaller in scale compared to high-resource languages. On the other hand, the quality of low-resource language corpora is often poor. The lack of standardized curation processes leads to numerous annotation errors and irregular records, which severely impact the accuracy of LLMs in understanding and processing these languages.
Model Capabilities Existing large language models have inherent limitations when dealing with low-resource languages. The model architectures are primarily designed based on extensive data from high-resource languages, making them less effective in handling the unique grammatical, lexical, and semantic features of low-resource languages. Additionally, the generalization capabilities of these models are insufficient. When encountering new vocabulary, expressions, or culturally specific semantics in low-resource languages, the models struggle to effectively utilize their training patterns for accurate understanding and generation.
Research Needs In humanities research, there are specific demands for processing low-resource languages. For instance, in ancient text translation and interpretation, models need to precisely understand the semantics and cultural connotations of ancient low-resource languages. In the preservation and study of minority cultures, models must comprehensively handle various texts involving minority languages. In the analysis of language evolution, models need deep insights into historical language changes. These demands require models to possess high precision and a deep understanding of linguistic nuances. However, there is a significant gap between current model capabilities and these research needs, particularly in handling complex metaphors, symbols, and other culturally loaded content in low-resource languages.
Limitations in Generative Technology Current LLM technology has some inherent limitations when dealing with low-resource languages. The pre-training data primarily consists of large-scale text collections dominated by high-resource languages, causing the models to overlook low-resource languages during initial learning. Consequently, when processing related tasks, the models lack prior knowledge, such as the ability to capture unique vocabulary usage, grammatical rules, and semantic networks in low-resource languages.
@@ 5.1.2 Challenges in Modeling Cultural Metaphors and Unique Usages
To address this challenge, LLMs need more robust contextual modeling and knowledge integration capabilities. On the one hand, building a more diverse and culturally representative training corpus helps models encounter more instances of cultural metaphors and special usages. On the other hand, incorporating external knowledge bases or knowledge graphs can provide the model with richer background information and reasoning abilities. Additionally, techniques such as contrastive learning and few-shot learning can enhance the model’s generalization capabilities in low-resource domains, enabling it to understand and generate these linguistic features even with limited examples. Ultimately, achieving culturally sensitive language modeling will open up new possibilities for cross-cultural communication and intelligent language interactions.
@@ 5.3.1 Opportunities: Role of Low-Resource Language Models
Low-resource language models have good potential in safeguarding endangered languages, traditions, and knowledge systems. These models empower the documentation and dissemination of oral traditions, folklore, and cultural narratives that might otherwise fade into obscurity due to a lack of balanced corpus resources. For example, the rich oral histories of indigenous communities, often passed down through generations, can be systematically transcribed and analyzed, ensuring their long-term preservation while simultaneously broadening their accessibility to a global audience. By bridging linguistic divides, language models[38] serve as conduits for cultural exchange, deepening cross-cultural intersectionality and fostering a sense of connection across diverse societies.
@@ 6.2 Identification of Specific Vocabulary and Concepts in Religious Texts
The LLMs provide revolutionary possibilities for the study of religious texts. In cross religious contexts, LLMs can compare and study traditional texts of different religions, identify terminology and doctrines. For example, LLMs can compare the concept of ”sympathy” expressed in Buddhist Pali scriptures, Christian New Testament texts, and Islamic hadith, providing insights into both general and specific traditional interpretations. LLMs can help explain the subtle differences of the same term in different religious backgrounds. For example, the Sanskrit term ”dharma” [8]in Hinduism may refer to justice or responsibility, while in Buddhism, it means the teachings of the Buddha or universal truth. LLMs trained extensively in context can help eliminate ambiguity in these terms during analysis or translation processes. In the cross-religious context, the same symbol may have different meanings in different traditions. For example, the lion is a symbol of strength and divinity in the Judeo Christian tradition, while in the Buddhist tradition it represents wisdom and courage. LLMs can compare these symbols in the corpus and enrich the interpretation by placing them in their respective traditions.
@@ 6.3.1 Opportunities: Innovative Applications of Language Models
The application of language models to low-resource literary and religious texts presents remarkable opportunities for preserving and revitalizing religious heritage. These models serve as vital bridges, reducing barriers to literary appreciation while making classic literary works more accessible to contemporary audiences. By delving into texts from low-resource languages such as Croatian and Quechua, language models can uncover and highlight the cultural nuances, historical context, and emotional depth embedded within these works, fostering deeper engagement and appreciation.
@@
In the realm of religious studies, language models offer a unique opportunity to address doctrinal conflicts and enrich theological discourse. By analyzing religious texts, these models can uncover principles in common and provide clarity in areas where doctrinal interpretations have historically diverged. For example, in Buddhism and Whiteness: Critical Reflections, the authors examine how entrenched racial and cultural dynamics shape religious interpretations, underscoring the importance to examine complex sociocultural factors within theological contexts. Furthermore, these models are capable of adapting and refining doctrinal interpretations to align with evolving perspectives[99], while remaining anchored in traditional frameworks. A notable application is the preservation and dissemination of religious traditions tied to specific cultural contexts, such as African-American theology[41]. With thoughtful development and ethical oversight, these models have the capacity to establish balanced frameworks that address doctrinal ambiguities while expanding the accessibility of religious knowledge across diverse cultural boundaries.
@@ 6.3.2 Challenges: Limitations in Models’ Understanding of Implicit Meaning
The limitation is compounded by the scarcity of preserved external resources and relevant data for many low-resource languages and traditions. While some texts have survived in written form, much of the oral tradition, cultural knowledge, and historical context that breathe life into these works have been lost to time. Without these necessary components, language models struggle to discern the deeper connections within the texts, frequently producing interpretations that are either reductive or outright erroneous. Meanwhile, LLMs often exhibit biases and cultural misrepresentations, as revealed by debate-induced evaluations, which highlight their tendency to align with dominant linguistic or cultural perspectives in training data[16], underscoring the need for fairness and contextual sensitivity in multilingual training.
@@ 7.2 Ethical Challenges
LLMs, while promising, can inadvertently worsen the marginalization of low-resource languages. Low-resource Indigenous languages often have limited digitized resources, leading to minimal and potentially flawed representation in language models. Uncorrected errors in machine-generated outputs risk amplification in subsequent models, creating a feedback loop that distorts authentic language usage  [98]. Furthermore, the dominance of high-resource languages in training corpora can lead to linguistic homogenization, further weakening the already vulnerable position of low-resource languages. This imbalance threatens the linguistic diversity essential for preserving unique cultural identities.

### v3

@@ 1.1.2 Demand for Low-Resource Languages in Humanities Research and the Insufficiencies in Existing Studies
The critical significance of low-resource languages notwithstanding, prevailing research methodologies encounter appreciable challenges when engaging with these languages. A principal limitation is the paucity of textual data. Numerous low-resource languages lack extensive written corpora, thereby complicating the execution of empirical studies and the development of computational models. Furthermore, the absence of specialized computational tools specifically adapted to these languages aggravates the issue [83]. Existing tools and resources are frequently designed for high-resource languages, thereby proving ineffective or inefficient for application to low-resource languages. As a brief empirical note from our working list, coverage remains uneven across LLM systems. For example, GPT5 supports Manchu whereas NLLB does not, while NLLB supports Istro-Venetian but GPT5 is not accurate. In short, each system covers some different languages, yet overall support for low-resource, including historical and regional, varieties remains insufficient for dependable humanities workflows. This technological disparity is further heightened by the inadequate availability of essential linguistic resources such as dictionaries, grammars, and annotated corpora, which are imperative for both descriptive and computational linguistic research.
@@ 4.1.2 Challenges in Modeling Cultural Metaphors and Unique Usages
To address this challenge, LLMs need more robust contextual modeling and knowledge integration capabilities. On the one hand, building a more diverse and culturally representative training corpus helps models encounter more instances of cultural metaphors and special usages. On the other hand, incorporating external knowledge bases or knowledge graphs can provide the model with richer background information and reasoning abilities. Additionally, techniques such as contrastive learning and few-shot learning can enhance the model’s generalization capabilities in low-resource domains, enabling it to understand and generate these linguistic features even with limited examples. Ultimately, developing culturally sensitive language modeling may improve cross-cultural communication and intelligent language interactions.
@@ 4.3.1 Opportunities: Role of Low-Resource Language Models
Low-resource language models may have good potential in safeguarding endangered languages, traditions, and knowledge systems. These models empower the documentation and dissemination of oral traditions, folklore, and cultural narratives that might otherwise fade into obscurity due to a lack of balanced corpus resources. For example, the rich oral histories of indigenous communities, often passed down through generations, can be systematically transcribed and analyzed, ensuring their long-term preservation while simultaneously broadening their accessibility to a global audience. By bridging linguistic divides, language models[59] serve as conduits for cultural exchange, deepening cross-cultural intersectionality and fostering a sense of connection across diverse societies.
@@ 5.2 Identification of Specific Vocabulary and Concepts in Religious Texts
Lai’s research[71] offers an interdisciplinary approach to linguistics, combining qualitative theory (Dialogic Syntax, Construction Grammar) with quantitative computational analysis focused on corpus linguistics, statistical modeling, and information theory. This ”usage-based” perspective views language as a dynamic system emerging from human interaction, exemplified by the analysis of texts like the Diamond Sutra. He argues that highly structured, non-spontaneous texts, like orally transmitted religious texts, provide a valuable, high-signal source for understanding linguistic patterns, suggesting a novel approach to LLM training by curating stylized, ritualized texts for enhanced structural and relational richness.
@@
The LLMs provide revolutionary possibilities[5] for the study of religious texts. In cross religious contexts, LLMs can compare and study traditional texts of different religions, identify terminology and doctrines. For example, LLMs can compare the concept of ”sympathy” expressed in Buddhist Pali scriptures, Christian New Testament texts, and Islamic hadith, providing insights into both general and specific traditional interpretations. LLMs can help explain the subtle differences of the same term in different religious backgrounds. For example, the Sanskrit term ”dharma” [15] in Hinduism may refer to justice or responsibility, while in Buddhism, it means the teachings of the Buddha or universal truth. LLMs trained extensively in context can help eliminate ambiguity in these terms during analysis or translation processes. In the cross-religious context, the same symbol may have different meanings in different traditions. Religious and philosophical texts are not simply repositories of information; they are highly structured performances of knowledge and authority, featuring unique rhetorical and structural patterns that are not present in everyday communication. A simple factual summary of such texts fails to capture their underlying rhetorical and dialogic function. For example, the lion is a symbol of strength and divinity in the Judeo Christian tradition, while in the Buddhist tradition it represents wisdom and courage. LLMs can compare these symbols in the corpus and enrich the interpretation by placing them in their respective traditions.
@@ 5.3.1 Opportunities: Innovative Applications of Language Models
The application of language models to low-resource literary and religious texts presents remarkable opportunities for preserving and revitalizing religious heritage. These models serve as vital bridges, reducing barriers to literary appreciation while making classic literary works more accessible to contemporary audiences. By delving into texts from low-resource languages such as Croatian and Quechua, language models can help identify cultural nuances, historical context, and emotional aspects embedded within these works, supporting broader engagement and deeper appreciation.
@@
Beyond accessibility and revitalization, LLMs can also be a powerful tool for academic and theological inquiry. In the realm of religious studies, language models offer a unique opportunity to address doctrinal conflicts and enrich theological discourse in several areas. By analyzing religious texts, these models can uncover common principles and provide clarity in areas where doctrinal interpretations have historically diverged. From a technological perspective, the systematically machine-learnable structure inherent in the Buddha-Subhti dialogue, as developed by Ryan Lai, presents several contributions to the fields of AI[70] and low-resource language processing. These contributions include the comprehensive modeling of epistemic status, the operationalization of rhetorical routines, and the recognition of the inherent value of stylized corpora for religious text, which furnish high-quality training data due to their structured textual nature. Through a profound understanding of linguistic routines and their associated social functions, this research endeavors to facilitate the development of more nuanced and communicatively proficient LLMs, particularly within intricate and specialized domains.
This article, alongside that of scholars like Carolyn Medine, underscores a crucial point: the successful application of LLMs in religious studies requires a comprehensive understanding of both the linguistic structure of sacred texts and their broader social and cultural contexts. For example, a purely textual analysis of Buddhism is insufficient without considering how entrenched racial and cultural dynamics shape its interpretations, as examined in [141]. Likewise, Carolyn Medine investigates the interdisciplinary culture of race, discrimination, and religion[97] in her tribute article to bell hooks, a black feminist and black buddhist, it highlights the need for LLMs to incorporate the nuances of race, gender, and social context in their analysis beyond a purely textual approach. Therefore, these models are capable of adapting and refining doctrinal interpretations to align with evolving perspectives, while remaining anchored in traditional frameworks. A notable application is the preservation and dissemination of religious traditions tied to specific cultural contexts, such as African-American theology[62]. With thoughtful development and ethical oversight, these models have the capacity to establish balanced frameworks that address doctrinal ambiguities while expanding the accessibility of religious knowledge across diverse cultural boundaries.
@@ 5.3.2 Challenges: Limitations in Models’ Understanding of Implicit Meaning
The limitation[94] is compounded by the scarcity of preserved external resources and relevant data for many low-resource languages and traditions. While some texts have survived in written form, much of the oral tradition, cultural knowledge, and historical context that breathe life into these works have been lost to time. In particular, religious and philosophical texts are highly structured performances of knowledge and authority, featuring unique rhetorical and structural patterns that are not present in everyday communication. A shallow factual summary of such texts fails to capture their underlying rhetorical and dialogic function. Without these necessary components, language models struggle to discern the deeper connections within the texts, frequently producing interpretations that are either reductive or outright erroneous. Meanwhile, LLMs often exhibit biases and cultural misrepresentations, as revealed by debate-induced evaluations, which highlight their tendency to align with dominant linguistic or cultural perspectives in training data[28], underscoring the need for fairness and contextual sensitivity in multilingual training.
@@ 6.2 Ethical Challenges
LLMs, while promising, can inadvertently worsen the marginalization of low-resource languages. Low-resource Indigenous languages often have limited digitized resources, leading to minimal and potentially flawed representation in language models. Uncorrected errors in machine-generated outputs risk amplification in subsequent models, creating a feedback loop that distorts authentic language usage  [138]. Furthermore, the dominance of high-resource languages in training corpora can lead to linguistic homogenization, further weakening the already vulnerable position of low-resource languages. This imbalance threatens the linguistic diversity and the preservation of cultural identities.